# Understanding New-Knowledge-Induced Factual Hallucinations in LLMs: Analysis, Solution and Interpretation

### v1

@@ Appendix B Training Details
In all CPT experiments, unless otherwise specified, we use a batch size of 16, a learning rate of 1e-5, a cutoff length of 512, and train for 1 epoch. In all SFT experiments (including knowledge QA and knowledge-based reasoning tasks), unless otherwise specified, we use a batch size of 32, a learning rate of 1e-5, and train for 3 epochs. We also did experiments of training 1 or 5 epochs, the results are presented in Appendix H. All experiments are conducted on up to 4 NVIDIA A6000 GPUs. The CPT stage is performed using LLaMA-Factory (Zheng et al. 2024).
@@
Appendix C Detailed Results of Main Text
@@
$\text{All}_{\text{k}}$ 0.549 0.609 0.546 0.464 0.199
$\text{B}_{\text{unk}}$ 0.200 0.591 0.555 0.466 0.195
$\text{D}_{\text{unk}}$ 0.539 0.225 0.531 0.464 0.188
$\text{M}_{\text{unk}}$ 0.545 0.596 0.255 0.456 0.183
$\text{U}_{\text{unk}}$ 0.531 0.606 0.546 0.252 0.157
@@ Appendix E CPT Results
The models are then evaluated with 5-shot QA format. We adapt the knowledge categorization method from Gekhman et al. (2024), with minor modifications. Specifically, we prompt the model with 5 different 5-shot samples. If the model answers correctly in at least one case, it is classified as Known; if all answers are incorrect, it is classified as Unknown. This is because the selection and order of few shots can significantly affect the model’s performance (Lu et al. 2022; Zhao et al. 2021), and we need to rule out this influence. The results are shown in Tables 10, 11, 12 and 13.
@@ Appendix F Supplementary Results of KnownPatch
Figure 10 is the result of KnownPatch on QA tasks with different injection ratios, in addition to Figure 4 in the main text; Figures 11 and 12 are results of KnownPatch on reasoning tasks with injection ratios of 5% and 10%, in addition to Figure 5 (results of injection ratio 5%) in the main text; Figures 13, 14 and 15 are detailed results of KnownPatch when one knowledge type is missed, with injection ratios 20%, 10% and 5%, respectively.
@@
Figure 13: KnownPatch (missing one knowledge type) on QA tasks with an injection ratio of 20%. All experiments trained for 3 epoch.
Figure 14: KnownPatch (missing one knowledge type) on QA tasks with an injection ratio of 10%. All experiments trained for 3 epoch.
Figure 15: KnownPatch (missing one knowledge type) on QA tasks with an injection ratio of 5%. All experiments trained for 3 epoch.
@@ Appendix G Results on Different Models
We used Qwen2.5-1.5B (main text) and Qwen3-8B, Llama3.2-1B (appendix), spanning architectures and sizes, all supporting our conclusions. Due to resource limitations, larger-scale training was infeasible. For larger models, prior work (Allen-Zhu and Li 2025) shows they retain factual knowledge better.
@@
Model B_QA D_QA M_QA U_QA wiki
$\text{All}_{\text{k}}$ 0.863 0.855 0.766 0.668 0.155
$\text{B}_{\text{unk}}$ 0.332 0.837 0.761 0.657 0.116
$\text{D}_{\text{unk}}$ 0.836 0.359 0.767 0.664 0.145
$\text{M}_{\text{unk}}$ 0.845 0.848 0.478 0.664 0.152
$\text{U}_{\text{unk}}$ 0.855 0.860 0.761 0.387 0.141
@@ G.2 Results on Qwen3-8B
Model B_QA D_QA M_QA U_QA wiki
$\text{All}_{\text{k}}$ 0.870 0.864 0.867 0.709 0.301
$\text{B}_{\text{unk}}$ 0.157 0.850 0.855 0.709 0.253
$\text{D}_{\text{unk}}$ 0.846 0.165 0.847 0.676 0.277
$\text{M}_{\text{unk}}$ 0.846 0.855 0.167 0.680 0.285
$\text{U}_{\text{unk}}$ 0.833 0.852 0.852 0.149 0.229
Table 15: Qwen3-8B model’s hallucination induced by training on different unknown knowledge types in QA tasks.
Figure 19: The impact of learning new knowledge in reasoning tasks on the average performance of different groups (on the Qwen3-8B model).
We perform the same analysis as Section 6 and Appendix D on the Qwen3-8B model. Based on the results of Figure 20 (similar to Figure 6), we chose its layers 9-27 for further interpretability analysis, and Figure 21 (similar to Figure 9) shows the results.
Figure 20: Qwen3-8B model’s attention score on the target name across layers in QA and reasoning training setups. The solid curves show the average attention score at each layer, aggregated across all datasets and instances. The shaded regions represent the standard deviation.
Figure 21: Qwen3-8B model’s performance and attention score changes when learning new knowledge in QA tasks, and after applying KnownPatch (with 20% known data). QA represents the average across the four QA test sets, and error bars indicate standard deviations.
@@ Appendix H Results on Different Training Epochs
Table 16 (similar to Table 2) provides the hallucination results in QA tasks when learning new knowledge; Figure 22 (similar to Figure 2) shows the performance after learning different proportions of unknown knowledge; Figure 23 (similar to Figure 3) shows the impact of new knowledge in reasoning tasks on different groups; Figure 24 (similar to Figure 10) reports the performance of KnownPatch on QA tasks when injecting 20% known data; Figure 25 (similar to Figure 5) reports performance of KnownPatch on reasoning tasks when injecting 20% known data; Figure 26 (similar to Figure 13) reports performance of KnownPatch when one knowledge type is missing with 20% injection ratio; Figure 27 (similar to Figure 8) reports the accuracy and attention score changes when learning new knowledge in reasoning tasks; Figure 28 (similar to Figure 7) reports the accuracy and attention score changes after learning different proportions of unknown knowledge; Figure 29 (similar to Figure 9) reports the performance and attention score changes before and after applying KnownPatch.
Model B_QA D_QA M_QA U_QA wiki
$\text{All}_{\text{k}}$ 0.548 0.568 0.475 0.458 0.198
$\text{B}_{\text{unk}}$ 0.198 0.527 0.481 0.448 0.199
$\text{D}_{\text{unk}}$ 0.518 0.280 0.471 0.442 0.190
$\text{M}_{\text{unk}}$ 0.517 0.576 0.193 0.423 0.173
$\text{U}_{\text{unk}}$ 0.529 0.594 0.467 0.304 0.174
@@
Figure 24: Performance of KnownPatch on QA task when injecting 20% known data. All experiments trained for 1 epoch.
@@
Table 17 (similar to Table 2) provides the hallucination results in QA tasks when learning new knowledge; Figure 30 (similar to Figure 2) shows the performance after learning different proportions of unknown knowledge; Figure 31 (similar to Figure 3) shows the impact of new knowledge in reasoning tasks on different groups; Figure 32 (similar to Figure 10) reports the performance of KnownPatch on QA task with 20% injection ratio; Figure 33 (similar to Figure 5) reports performance of KnownPatch when injecting 20% known data; Figure 34 (similar to Figure 13) reports performance of KnownPatch when one knowledge type is missing when injecting 20% known data; Figure 35 (similar to Figure 8) reports the accuracy and attention score changes when learning new knowledge in reasoning tasks; Figure 36 (similar to Figure 7) reports the accuracy and attention score changes after learning different proportions of unknown knowledge; Figure 37 (similar to Figure 9) reports the performance and attention score changes before and after applying KnownPatch.
Model B_QA D_QA M_QA U_QA wiki
$\text{All}_{\text{k}}$ 0.529 0.600 0.552 0.447 0.200
$\text{B}_{\text{unk}}$ 0.234 0.573 0.550 0.434 0.194
$\text{D}_{\text{unk}}$ 0.505 0.289 0.533 0.435 0.181
$\text{M}_{\text{unk}}$ 0.523 0.606 0.215 0.432 0.171
$\text{U}_{\text{unk}}$ 0.523 0.607 0.554 0.245 0.144
@@
Figure 32: Performance of KnownPatch on QA task when injecting 5% (upper), 10% (middle) and 20% (lower) known data. All experiments trained for 5 epoch.
@@
Table 18 (similar to Table 2) provides the hallucination results in QA tasks when learning new knowledge; Figure 38 (similar to Figure 2) shows the performance after learning different proportions of unknown knowledge; Figure 39 (similar to Figure 3) shows the impact of new knowledge in reasoning tasks on different groups; Figure 40 (similar to Figure 10) reports the performance of KnownPatch on QA task when injecting 20% known data; Figure 41 (similar to Figure 5) reports performance of KnownPatch when injecting 20% known data; Figure 42 (similar to Figure 13) reports performance of KnownPatch when one knowledge type is missing when injecting 20% known data; Figure 43 (similar to Figure 8) reports the accuracy and attention score changes when learning new knowledge in reasoning tasks; Figure 44 (similar to Figure 7) reports the accuracy and attention score changes after learning different proportions of unknown knowledge; Figure 45 (similar to Figure 9) reports the performance and attention score changes before and after applying KnownPatch.
Model B_QA D_QA M_QA U_QA wiki
$\text{All}_{\text{k}}$ 0.498 0.611 0.540 0.371 0.186
$\text{B}_{\text{unk}}$ 0.183 0.503 0.516 0.380 0.176
$\text{D}_{\text{unk}}$ 0.443 0.231 0.499 0.352 0.159
$\text{M}_{\text{unk}}$ 0.469 0.544 0.343 0.360 0.178
$\text{U}_{\text{unk}}$ 0.452 0.549 0.485 0.252 0.171
@@
Figure 40: Performance of KnownPatch on QA task when injecting 5% (upper), 10% (middle) and 20% (lower) known data. All experiments trained for 20 epoch.

### v2

@@ Appendix B Training Details
In all CPT experiments, unless otherwise specified, we use a batch size of 16, a learning rate of 1e-5, a cutoff length of 512, and train for 1 epoch.
In all SFT experiments (including knowledge QA and knowledge-based reasoning tasks), unless otherwise specified, we use a batch size of 32, a learning rate of 1e-5, and train for 3 epochs. We also did experiments of training 1 or 5 epochs, the results are presented in Appendix H.
In the experiment that uses an auxiliary KL loss in Section 5.2, due to this additional constraint, training becomes more difficult, so we set the number of training epochs to at most ten and apply early stopping when the training accuracy exceeds 95%. All other training setups are the same as above.
Most of the experiments are conducted on up to four NVIDIA A6000 GPUs for models with fewer than 8B parameters. For training models with 8B parameters or more, up to eight NVIDIA H20 GPUs are used. The CPT stage is performed using LLaMA-Factory (Zheng et al., 2024).
@@
Appendix C More Results for Main Text
C.1 Detailed Results for Main Text
@@
$\text{All}_{\text{k}}$ 0.549 0.650 0.519 0.442 0.199
$\text{B}_{\text{unk}}$ 0.252 0.618 0.513 0.430 0.179
$\text{D}_{\text{unk}}$ 0.521 0.322 0.511 0.426 0.168
$\text{M}_{\text{unk}}$ 0.539 0.627 0.244 0.445 0.165
$\text{U}_{\text{unk}}$ 0.528 0.635 0.508 0.179 0.143
@@
C.2 Results on Real-World Dataset
To examine whether our findings hold on real-world data, we additionally conduct experiments on the ENTITYQUESTIONS dataset (Sciavolino et al., 2021). A key difference between real-world data and our synthetic dataset lies in whether the corresponding knowledge has already been observed by the model during its initial pre-training stage.
Following the procedure described in Appendix A.1, we categorize knowledge for the Qwen2.5-1.5B model into four groups: HighlyKnown, MaybeKnown, WeaklyKnown, and Unknown. We select four knowledge types that contain a relatively large proportion of HighlyKnown and Unknown instances: P36, P106, P159, and P407, and use them as training knowledge types.
Following the experimental setups in Section 4.1, we sample 500 instances for each knowledge type. A model trained on the fully known dataset serves as the baseline model, while variant models are obtained by replacing one knowledge type with unknown instances. We then evaluate these models on the corresponding test sets of the four knowledge types, as well as on the same Wiki test set used in the main text. The results are shown in Table 13, which shows exactly the same trends as Table 2.
STQA DTQA wiki
-36.94 ($\pm$ 11.30) -3.29 ($\pm$ 3.35) -6.86 ($\pm$ 2.57)
Table 13: Average performance degradation (%, mean ± std) of four model variants.
We conduct experiments to examine how the proportion of unknown knowledge affects the severity of hallucinations. The results are shown in Figure 8, which shows similar trends as Figure 2.
Figure 8: Performance with different proportions of unknown knowledge in the same type and Wiki test set.
Following the experiments in Section 5.2, in Figure 9 we also show that the attention to key entities is correlated to performance drop when unknown knowledge proportion increases.
Figure 10 presents the results of KnownPatch (with injection ratio 20%) on real-world data, along with its attention analysis.
Figure 10: Performance and attention score changes under Shuffled and KnownPatch (with 20% injection ratio) settings. QA represents the average across the four QA test sets, and error bars indicate standard deviations.
@@ Appendix E CPT Results
The models are then evaluated with 5-shot QA format. We adapt the knowledge categorization method from Gekhman et al. (2024), with minor modifications. Specifically, we prompt the model with 5 different 5-shot samples. If the model answers correctly in at least one case, it is classified as Known; if all answers are incorrect, it is classified as Unknown. This is because the selection and order of few shots can significantly affect the model’s performance (Lu et al., 2022; Zhao et al., 2021), and we need to rule out this influence. The results are shown in Tables 14, 15, 16 and 17.
@@ Appendix F Supplementary Results of KnownPatch
Figure 11 is the result of KnownPatch on QA tasks with different injection ratios; Figures 12, 13 and 14 are results of KnownPatch on reasoning tasks with injection ratios of 5%, 10% and 20%; Figures 15 and 16 are interpretability results of injecting 10% and 5% known data in KnownPatch.
@@
Figure 14: KnownPatch on reasoning tasks with 20% injection ratio. All experiments trained for 3 epoch.
Figure 15: Performance and attention score changes when learning new knowledge in QA tasks, and after applying KnownPatch (with 10% known data). QA represents the average across the four QA test sets, and error bars indicate standard deviations.
Figure 16: Performance and attention score changes when learning new knowledge in QA tasks, and after applying KnownPatch (with 5% known data). QA represents the average across the four QA test sets, and error bars indicate standard deviations.
@@ Appendix G Results on Different Models
We used Qwen2.5-1.5B (main text) and Qwen3-8B, Llama3.2-1B (appendix), spanning architectures and sizes, all supporting our conclusions. Due to resource limitations, larger-scale training was infeasible. For larger models, prior work (Allen-Zhu and Li, 2025) shows they retain factual knowledge better.
@@
STQA DTQA Wiki
-49.80 ($\pm$ 11.74) -1.04 ($\pm$ 1.03) -10.65 ($\pm$ 10.11)
@@ G.2 Results on Qwen3-8B-Base
STQA DTQA Wiki
-80.64 ($\pm$ 1.23) -2.33 ($\pm$ 1.42) -13.29 ($\pm$ 8.40)
Table 19: Qwen3-8B-Base model’s hallucination induced by training on different unknown knowledge types in QA tasks.
Figure 20: The impact of learning new knowledge in reasoning tasks on the average performance of different groups (on the Qwen3-8B-Base model).
We perform the same analysis as Section 5 and Appendix D on the Qwen3-8B-Base model. Based on the results of Figure 21 (similar to Figure 4), we chose its layers 9-27 for further interpretability analysis, and Figure 22 (similar to Figure 7) shows the results.
Figure 21: Qwen3-8B-Base model’s attention score on the target name across layers in QA and reasoning training setups. The solid curves show the average attention score at each layer, aggregated across all datasets and instances. The shaded regions represent the standard deviation.
Figure 22: Qwen3-8B-Base model’s performance and attention score changes when learning new knowledge in QA tasks, and after applying KnownPatch (with 20% known data). QA represents the average across the four QA test sets, and error bars indicate standard deviations.
G.3 Results on Qwen2.5-32B
STQA DTQA Wiki
-74.00 ($\pm$ 5.29) -2.59 ($\pm$ 3.57) -9.69 ($\pm$ 7.88)
Table 20: Qwen2.5-32B model’s hallucination induced by training on different unknown knowledge types in QA tasks.
Figure 23: The impact of learning new knowledge in reasoning tasks on the average performance of different groups (on the Qwen2.5-32B model).
We perform the same analysis as Section 5 and Appendix D on the Qwen2.5-32B model. Based on the results of Figure 24 (similar to Figure 4), we chose its layers 25-55 for further interpretability analysis, and Figure 25 (similar to Figure 7) shows the results.
Figure 24: Qwen2.5-32B model’s attention score on the target name across layers in QA and reasoning training setups. The solid curves show the average attention score at each layer, aggregated across all datasets and instances. The shaded regions represent the standard deviation.
Figure 25: Qwen2.5-32B model’s performance and attention score changes when learning new knowledge in QA tasks, and after applying KnownPatch (with 20% known data). QA represents the average across the four QA test sets, and error bars indicate standard deviations.
@@ Appendix H Results on Different Training Epochs
Table 21 (similar to Table 2) provides the hallucination results in QA tasks when learning new knowledge; Figure 26 (similar to Figure 2) shows the performance after learning different proportions of unknown knowledge; Figure 27 (similar to Figure 3) shows the impact of new knowledge in reasoning tasks on different groups; Figure 28 (similar to Figure 14) reports performance of KnownPatch on reasoning tasks when injecting 20% known data; Figure 30 (similar to Figure 6) reports the accuracy and attention score changes when learning new knowledge in reasoning tasks; Figure 31 (similar to Figure 5) reports the accuracy and attention score changes after learning different proportions of unknown knowledge; Figure 32 (similar to Figure 7) reports the performance and attention score changes before and after applying KnownPatch.
STQA DTQA Wiki
-51.89 ($\pm$ 13.35) -2.53 ($\pm$ 3.72) -7.07 ($\pm$ 6.40)
@@
Table 22 (similar to Table 2) provides the hallucination results in QA tasks when learning new knowledge; Figure 33 (similar to Figure 2) shows the performance after learning different proportions of unknown knowledge; Figure 34 (similar to Figure 3) shows the impact of new knowledge in reasoning tasks on different groups; Figure 35 (similar to Figure 14) reports performance of KnownPatch when injecting 20% known data; Figure 37 (similar to Figure 6) reports the accuracy and attention score changes when learning new knowledge in reasoning tasks; Figure 38 (similar to Figure 5) reports the accuracy and attention score changes after learning different proportions of unknown knowledge; Figure 39 (similar to Figure 7) reports the performance and attention score changes before and after applying KnownPatch.
STQA DTQA Wiki
-53.46 ($\pm$ 6.68) -1.79 ($\pm$ 2.05) -13.75 ($\pm$ 10.60)
@@
Table 23 (similar to Table 2) provides the hallucination results in QA tasks when learning new knowledge; Figure 40 (similar to Figure 2) shows the performance after learning different proportions of unknown knowledge; Figure 41 (similar to Figure 3) shows the impact of new knowledge in reasoning tasks on different groups; Figure 42 (similar to Figure 14) reports performance of KnownPatch when injecting 20% known data; Figure 44 (similar to Figure 6) reports the accuracy and attention score changes when learning new knowledge in reasoning tasks; Figure 45 (similar to Figure 5) reports the accuracy and attention score changes after learning different proportions of unknown knowledge; Figure 46 (similar to Figure 7) reports the performance and attention score changes before and after applying KnownPatch.
STQA DTQA Wiki
-48.50 ($\pm$ 16.53) -7.73 ($\pm$ 5.06) -8.06 ($\pm$ 4.58)